why is the answer to the summation notation for the series of numbers {2,4,6,...100} equal 50x51 ?
Answer #1

The sum of first n natural even numbers is given by n*(n+1) where n is total no.

Here, we are given a series of even numbers from 2 to 100, thus there are a total of 50 numbers.

Thus its sum is given by n*(n+1) = 50*51= 2550